How do you solve the following problems?
(Im not able to use the radicals sign, so try to solve it imagining its there, please)
rs=radical sign
3 rs24

Answer by RAY100(1637)

3sqrt24=3*sqrt(4*6) = 3{sqrt4*sqrt6}= 3*2sqrt6=6sqrt6
.
check, use calc on original and answer,,,both = 14.7,,,ok
.